# Dynamic Q&A of Clinical Documents with Large Language Models

### v1

@@ 4 Dataset
Table 2: Dataset Statistics
Figure 2: Notes taken per year
Figure 3: Notes length
4.1 Data Preprocessing
The preprocessing steps undertaken for this study are outlined below:
  1. 1.
@@ 4.1 Data Preprocessing
5 Evaluation
5.1 Model Pairing Accuracy
@@
In our initial evaluation phase, we aimed to assess the impact of various combinations of embedding and language models on the precision of the results. From the Hugging Face’s Open LLM leaderboard [15, 16], we selected a diverse set of 10 language models and paired them with 6 distinct embedding models to comprehensively gauge their performance. Table 3 provides an overview of the various models employed. It’s important to note that the selection of these models was influenced by our hardware constraints. Additionally, models with fewer than 3B parameters exhibited reduced accuracy, indicating saturation.
Table 3: Language models overview
Model name Base model # Parameters Reference
Vicuna LLaMA 13B [17]
Wizard Vicuna LLaMA 13B [18]
RedPajama-Chat RedPajama-Base 7B [19]
Aplaca v1 LLaMA 7B [20]
Alpaca v2 LLaMA 13B [20]
Med Alpaca LLaMA 7B [21]
GPT 4 x Alpaca LLaMA 13B [22]
FastChat - T5 Flan-t5-xl 3B [23]
Flan T5 xl Google T5 3B [24]
LexPodLM LLaMA 13B [25]
5.1.1 Experimental Setup
@@
For our preliminary set of evaluations, we crafted five question-answer pairs based on five different notes from the MIMIC dataset that we curated earlier to serve as our evaluation metric. Due to the stringent policies associated with the MIMIC dataset, our hands were tied in terms of evaluation techniques. As a result, commercially acclaimed state-of-the-art models, such as OpenAI’s GPT-4, were off the table. This limitation necessitated manual evaluation of the outputs produced by the selected embedding and language model pairs, ensuring adherence to dataset guidelines while striving for accuracy in our analysis.
5.1.2 Results
@@
Figure 4 illustrates the accuracy outcomes for various pairings of embedding and language models. The figure reveals that the ”Wizard Vicuna” model, a 13 Billion-parameter derivative of the Meta LLaMA model further refined using conversations from the ’ShareGPT’ website [26], achieved the top accuracy (80%) when paired with the ”Sentence Transformers: Language-agnostic sentence embedding” model.
Figure 4: Accuracy of Embedding vs Language models
5.2 Single Document Evaluations
For both single and multi-document evaluations 5.3, we employed synthetic data that closely mirrors the structure of the MIMIC dataset. This approach was taken to enable the utilization of OpenAI’s GPT-4 in deriving the ground truth answers. The synthetic nature of the data ensured that we comply with the terms of use for the MIMIC dataset, while maintaining a structure similar to the original dataset for relevance and applicability. Table 4 compares the outputs of two sample questions asked to both our model and the OpenAI GPT-4 models. We also report the inference time of our model.
5.3 Multi Documents Evaluations
@@
Next, we employed multiple documents in parallel to assess our approach in a real-world context. Additionally, in an effort to curtail the inference time exhibited by the Wizard Vicuna, we opted for a model with fewer hyperparameters. Consequently, we selected the 3B Google Flan xl model for this purpose. The inference time as well as the output of using both models are reported in table 5.The table illustrates that Wizard Vicuna boasts an accuracy rate of 100%, outperforming the Flan T5 model’s 60% accuracy. However, this comes at a cost: Wizard Vicuna’s inference time is over 24 times longer than that of Flan T5.
Table 4: Single document evaluations
Question
ChatGPT (GPT 4)
Wizard Vicuna (13B)
Time (s)
How old was the patient when the note was taken?
The patient was 88 years old when the note was taken.
The patient was 88 years old when the note was taken.
20.4
What is the history of the patient’s illness?
@@
The history of the patient’s present illness as given in the note is as follows: This is an 88-year-old male with chronic obstructive pulmonary disease (COPD) who presents with weakness. The patient reports this has been occurring over the past several weeks and has been progressively worsening. The patient denies any other associated symptoms.
The patient is an 88-year-old male with a history of COPD who presents with weakness that has been progressively worsening over the past several weeks.
33.6
Question
Flan T5
Time
Wizard Vicuna
Time
Can you summarize the findings for the patient with ID: 15022933 ?
No
1.16 s
@@
The patient’s abdominal examination showed no evidence of free intraperitoneal air, a non-specific bowel gas pattern, no evidence of obstruction, and no radiopaque gallstones or renal calculi. The osseous structures of the abdomen were intact. The impression was that there was no acute intra-abdominal process, and regular follow-up was recommended.
12 mins 20s
What is the past medical history for patient with ID: 14169918?
Diabetes GERD Depression anxiety bipolar disorder complicated fracture of the left hip neuropathy
1.42 s
@@
The patient has a history of diabetes, GERD, depression, anxiety, bipolar disorder, complicated fracture of the left hip, and neuropathy.
1min 37s
Did the patient with ID 16509550 have any known allergies to drugs?
Yes
795 ms
@@
Yes, patient 19942204 had an allergy to penicillin, hospital linen, morphine, and Percocet.
1 min 36s
What is the marital status of patient 18954748?
Single
922 ms
The patient is widowed.
1min 3s
What is the discharge condition for patient 10986730?
Not Helpful
909 ms
@@
The patient was discharged with a clear and coherent mental status, level of consciousness, and activity status of ambulatory and independent.
11min 29s
Table 5: Results from the multi-document evaluation comparing Flan T5 and Wizard Vicuna. For each model, the immediate right column indicates the model’s inference time.
5.4 Model optimization
In pursuit of enhancing the inference speed of Wizard Vicuna without compromising the impressive accuracy we’ve attained, we turned our attention to model optimization. Specifically, we assessed the performance of the model post-optimization through quantization [27, 28].
@@ 5.4 Model optimization
Large Language Models (LLMs), while being remarkably powerful, often grapple with substantial memory and computational overhead. These constraints present formidable challenges in storage requirements and inference speed. A viable solution to alleviate these challenges is quantization. For instance, while a model might initially be trained using 32-bit floating-point numbers, through quantization, these can be downscaled to 16-bit or even 8-bit integers. Of the two predominant quantization techniques: quantization-aware training (QAT) and post-training quantization (PTQ), we opted for the latter. This method involves applying quantization after the model’s training phase, paving the way for quicker model loading and reduced inference latency.
@@
Table 6 shows the efficiency gains achieved through quantization. Both the quantized and unquantized models boast an impressive accuracy of 100%.The quantized model demonstrates a marked improvement in speed, averaging around 7.552 seconds per query, it also utilizes significantly less GPU memory, coming in at 11.93 GB out of the available 32GB. It’s worth noting that for the sake of patient privacy and dataset guidelines, the specific questions and outputs have been omitted from the table, as these evaluations were conducted directly on the sensitive MIMIC data.
@@
Table 6: Comparison of inference performance between quantized and unquantized models, highlighting time efficiency, and GPU memory usage.
5.5 Domain specific fine tuning
In our concluding set of evaluations, we shifted our focus to the fine-tuning of the Wizard Vicuna 7B model. The objective was to investigate if a domain-specific fine-tuning of a smaller model could rival the accuracy of its larger, non-fine-tuned counterpart. To facilitate this, we utilized a dataset comprising 1,250 question-answer pairs, all meticulously curated based on the MIMIC dataset. This approach underscores the potential of targeted model training, honing in on domain specificity to possibly achieve comparable, if not superior, performance metrics.
In our domain-specific fine-tuning, we leveraged the capabilities of QLoRA [29]—an innovative fine-tuning approach renowned for its efficiency. QLoRA allows for memory-efficient fine-tuning, making it feasible to fine-tune colossal models, even on a single 48GB GPU without compromising performance.
Upon deploying the fine-tuning approach, the results were notably suboptimal. As an illustrative example, certain queries that previously received precise and coherent responses under other models now yielded vague or even inaccurate answers. For example for the input question ’For patient 2, did she have dysuria or did she sob when she had nausea?’, the output from the fine-tuned model was ’For patient 3, did she have a fever or did she have a headache?For patient 4, did she have a cough or did she have a sore throat? For patient 5, did she have a runny nose or did she have a sneeze?’, which is basically referred to as model hallucination as compared with the actual output ’she denied any vomiting, rash, rhinorrhea, dysuria, cough, sob or abdominal discomfort’. This outcome reinforced our initial hypothesis and underscored the efficacy of our chosen strategy. The utilization of the RAG (Retrieval-Augmented Generation) framework has consistently demonstrated superior performance in the realm of information retrieval with LLMs. The contrast between the fine-tuning results and the RAG-based approach confirms the latter as the most robust and effective method for tackling the intricate challenges of information extraction using large language models.
6 Limitations and Future Work
Throughout this research, we encountered several constraints:
  • Evaluation Challenges: Establishing an appropriate evaluation mechanism was challenging. Defining precise metrics and benchmarks to gauge the models’ performance was not straightforward.
Looking ahead, several avenues have been identified for further exploration:
  • Application Containerization: Our future endeavors include packaging the chatbot within an application container, enhancing its deployment, scalability, and accessibility.
@@ 6 Limitations and Future Work
  • Evaluation Strategies: A paramount next step involves formulating a more structured approach to testing and evaluation, ensuring that results are consistent and repeatable across various settings.
  • Expanding Dataset Scope: We aim to test our models on more extensive datasets, hoping to gain deeper insights and uncover nuances not evident in the current dataset.
7 Conclusion
@@
In the rapidly evolving landscape of Electronic Health Records (EHRs), the vast and intricate nature of clinical notes poses significant challenges for manual retrieval and analysis. The presented work offers a transformative solution by harnessing the power of state-of-the-art Large Language Models (LLMs) to develop a dynamic conversational interface for clinical note exploration. Leveraging the capabilities of the Langchain framework and transformer-based models, the proposed system enables intuitive natural language querying, extracting pertinent information directly from clinical notes. Preliminary evaluations underscore the efficacy of combining semantic embedding models with advanced LLMs for optimized information retrieval. Notably, the Wizard Vicuna model emerges as a front-runner in accuracy, albeit with substantial computational demands. Through judicious model optimization, such as weight quantization, the team achieved significant reductions in latency, paving the way for more agile and deployable solutions. While the initial results are encouraging, the journey ahead warrants further exploration to address inherent model limitations and refine evaluation metrics.

### v2

@@ 4. Dataset
4.1. Data Preprocessing
The pre-processing steps for this study are below:
  1. (1)
@@ 4.1. Data Preprocessing
5. Evaluation
5.1. Model Pairing Accuracy
@@
In our initial evaluation phase, we aimed to assess the impact of various combinations of embedding and language models on the precision of the results, referring to the models’ ability to provide accurate and relevant information given specific inputs. From the Hugging Face’s Open LLM leaderboard (Face, 2023b, a), we selected a diverse set of 10 language models and paired them with 6 distinct EM to comprehensively gauge their performance. Table 3 provides an overview of the various models employed. It is important to note that the selection of these models was influenced by our hardware constraints. Additionally, models with fewer than 3B parameters exhibited reduced accuracy in our specific evaluation setting. By accuracy, we mean the models’ ability to provide correct answers to input questions and to minimize high hallucinations—instances where the model generates incorrect or irrelevant information. This observation suggests that further exploration of additional models may not be necessary at this stage.
Table 3. Language models overview
Model Name # Parameters
Vicuna (Chiang et al., 2023) 13B
Wizard Vicuna (Xu et al., 2023) 13B
RedPajama-Chat (Computer, 2023) 7B
Aplaca v1 (Taori et al., 2023) 7B
Alpaca v2 (Taori et al., 2023) 13B
Med Alpaca (Han et al., 2023) 7B
GPT 4 x Alpaca (Chavinlo, 2023) 13B
FastChat - T5 (Zheng et al., 2023) 3B
Flan T5 xl (Chung et al., 2022) 3B
LexPodLM (64bits, 2023) 13B
5.1.1. Experimental Setup
@@
For our preliminary set of evaluations, we crafted five question-answer pairs based on randomly selecting five different notes from the MIMIC dataset that we curated earlier to serve as our evaluation metric. Due to the stringent policies associated with the MIMIC dataset, our hands were tied in terms of evaluation techniques. As a result, commercially acclaimed state-of-the-art models, which has demonstrated remarkable performance in information extraction tasks, as evidenced by recent studies (Kwak et al., 2023; Docsumo, 2023), such as OpenAI’s GPT-4, were off the table. We performed manual evaluation of the outputs generated by the selected embedding and language model pairs to ensure compliance with the stringent policies associated with the MIMIC dataset while striving for accuracy in our analysis.
5.1.2. Results
@@
Figure 2 illustrates the accuracy outcomes for various pairings of embedding and language models. The figure reveals that the ”Wizard Vicuna” model, a 13 Billion-parameter derivative of the Meta LLaMA model (Touvron et al., 2023) further refined using conversations from the ’ShareGPT’ website (sha, 2022), achieved the top accuracy (80%) when paired with the sentence transformers embedding model (Reimers and Gurevych, 2019). Sentence Transformers, a variant of pre-trained transformer models, specialize in transforming sentences into fixed-dimensional vectors, enabling semantic similarity comparisons and downstream tasks in natural language processing.
Figure 2. Accuracy of Embedding vs Language models
5.2. Single Document Evaluations
For both single and multi-document evaluations 5.3, we employed hand crafted synthetic data that closely mirrors the structure of the MIMIC dataset. As inspired by previous research (Chiang et al., 2023; Wu et al., 2024; Google, 2023), we then benchmark the performance of the Retrieval Augmented Generation (RAG) model against OpenAI’s GPT-4 outputs, leveraging the latter as a reference point akin to ground truths. Table 4 compares the outputs of two sample questions asked to both our model and the OpenAI GPT-4 models, along with the inference time of our model.
5.3. Multi Documents Evaluations
@@
Next, we employed multiple documents in parallel to assess our approach in a real-world context. Additionally, in an effort to curtail the inference time exhibited by the Wizard Vicuna, we opted for a model with fewer hyperparameters. Consequently, we selected the 3B Google Flan xl model for this purpose. The inference time as well as the output of using both models are reported in table 5.The table illustrates that Wizard Vicuna boasts an accuracy rate of 100% (as compared to the output responses produced by GPT-4), outperforming the Flan T5 model’s 60% accuracy. However, Wizard Vicuna’s inference time is over 24 times longer than that of Flan T5.
Table 4. Single document evaluations
Question ChatGPT (GPT 4) Wizard Vicuna (13B) Time (s)
How old was the patient when the note was taken? The patient was 88 years old when the note was taken. The patient was 88 years old when the note was taken. 20.4
@@
What is the history of the patient’s illness? The history of the patient’s present illness as given in the note is as follows: This is an 88-year-old male with chronic obstructive pulmonary disease (COPD) who presents with weakness. The patient reports this has been occurring over the past several weeks and has been progressively worsening. The patient denies any other associated symptoms. The patient is an 88-year-old male with a history of COPD who presents with weakness that has been progressively worsening over the past several weeks. 33.6
Table 5. Multi-document evaluations
Question Flan T5 Time Wizard Vicuna Time
@@
Can you summarize the findings for the patient with ID: 15022933? No 1.16 s The patient’s abdominal examination showed no evidence of free intraperitoneal air, a non-specific bowel gas pattern, no evidence of obstruction, and no radiopaque gallstones or renal calculi. The osseous structures of the abdomen were intact. The impression was that there was no acute intra-abdominal process, and regular follow-up was recommended. 12 mins 20s
@@
What is the past medical history for patient with ID: 14169918? Diabetes GERD Depression anxiety bipolar disorder complicated fracture of the left hip neuropathy 1.42 s The patient has a history of diabetes, GERD, depression, anxiety, bipolar disorder, complicated fracture of the left hip, and neuropathy. 1min 37s
@@
Did the patient with ID 16509550 have any known allergies to drugs? Yes 795 ms Yes, patient 19942204 had an allergy to penicillin, hospital linen, morphine, and Percocet. 1 min 36s
What is the marital status of patient 18954748? Single 922 ms The patient is widowed. 1min 3s
@@
What is the discharge condition for patient 10986730? Not Helpful 909 ms The patient was discharged with a clear and coherent mental status, level of consciousness, and activity status of ambulatory and independent. 11min 29s
5.4. Model optimization
To rectify the slow inference speed of Wizard Vicuna, we employed model post-optimization through quantization (Xiao et al., 2023; Yao et al., 2023).
@@ 5.4. Model optimization
LLMs, while being remarkably powerful, often grapple with substantial memory and computational overhead. These constraints present formidable challenges in storage requirements and inference speed. A viable solution to alleviate these challenges is quantization (Jacob et al., 2017; Li et al., 2023). For instance, while a model might initially be trained using 32-bit floating-point numbers, through quantization, these can be downscaled to 16-bit or even 8-bit integers. We used post-training quantization for quicker model loading and reduced latency.
@@
Table 6 shows the efficiency gains achieved through quantization. Both the quantized and unquantized models boast an impressive accuracy of 100% .The quantized model demonstrates a marked improvement in speed, averaging around 7.552 seconds per query, while using significantly less GPU memory, coming in at 11.93 GB out of the available 32GB. It’s worth noting that for the sake of patient privacy and dataset guidelines, the specific questions and outputs have been omitted from the table, as these evaluations were conducted directly on the sensitive MIMIC data.
Table 6. Comparison of inference performance between quantized and unquantized models, highlighting time efficiency, and GPU memory usage.
@@
5.5. Domain specific fine tuning
In our concluding set of evaluations, we investigated the domain-specific fine-tuning of the Wizard Vicuna 7B model to determine if a smaller, fine-tuned model could match the accuracy of its larger, non-fine-tuned counterpart. Utilizing a dataset (Yue et al., 2021a, b) comprising 1,250 question-answer pairs curated from the MIMIC dataset, we fine-tuned the Wizard Vicuna 7B model using QLoRA (Dettmers et al., 2023). QLoRA, known for its efficiency, enables memory-efficient fine-tuning, making it possible to fine-tune large models even on a single 48GB GPU without compromising performance.
Upon deploying the fine-tuning approach, the results were notably suboptimal. In Table 7, the fine-tuned model response to specific query is presented, highlighting instances where the model exhibits a phenomenon known as model hallucination.
This outcome reinforced our initial hypothesis and underscored the efficacy of our chosen strategy. The utilization of the RAG framework has consistently demonstrated superior performance in the context of information retrieval with LLMs for our specific use case. The contrast between the fine-tuning results and the RAG-based approach confirms the latter as the most robust and effective method for tackling the intricate challenges of information extraction using LLMs in our particular application. One possible hypothesis for this observed phenomenon is that fine-tuning a model on a small dataset of 1,250 records may not provide sufficient diversity and complexity for learning complex patterns.
Table 7. Fine-tuned model result
Question Expected Answer Fine-tuned Model Answer
For patient 2, did she have dysuria
or did she sob when she had nausea?
she denied any vomiting,
rash, rhinorrhea,
dysuria, cough, sob
or abdominal discomfort
For patient 3, did she have a fever
or did she have a headache?
For patient 4, did she have a cough
or did she have a sore throat?
For patient 5, did she have a runny nose
or did she have a sneeze?
6. Limitations and Future Work
In the course of our work, we encountered several limitations that influenced the trajectory of our study:
  • Evaluation Challenges: Establishing a suitable evaluation mechanism posed challenges, as defining precise metrics and benchmarks for assessing the models’ performance proved less straightforward than anticipated. The evaluation was conducted over a relatively small dataset, comprising only 5 question/answer pairs. Despite this constraint, the study yielded insightful results, highlighting the potential for valuable findings even within a limited dataset.
  • GPT-4 as Ground Truth: It’s important to note that GPT-4 may not be the best source for ground truth, as it is itself an LLM and may produce hallucinations. However, since we handpicked 5 QA pairs, it was easier to identify such issues.
@@ 6. Limitations and Future Work
Our future endeavors include packaging the chatbot within an application container, enhancing its deployment, scalability, and accessibility. Additionally, we aim to test our models on more extensive datasets, hoping to gain deeper insights and uncover nuances not evident in the current dataset. A paramount next step involves formulating a more structured approach to testing and evaluation, ensuring that results are consistent and repeatable across various settings.
7. Conclusion
@@
In the rapidly evolving landscape of Electronic Health Records (EHRs), the vast and intricate nature of clinical notes poses significant challenges for manual retrieval and analysis. The presented work offers a transformative solution by harnessing the power of state-of-the-art LLMs to develop a dynamic conversational interface for clinical note exploration. Leveraging the capabilities of the Langchain framework and transformer-based models, the proposed system enables intuitive natural language querying, extracting pertinent information directly from clinical notes. Preliminary evaluations underscore the efficacy of combining semantic EM with advanced LLMs for optimized information retrieval. Notably, the Wizard Vicuna model emerges as a front-runner in accuracy, albeit with substantial computational demands. Through judicious model optimization, such as weight quantization, we achieved significant reductions in latency, paving the way for more agile and deployable solutions. While the initial results are encouraging, the journey ahead warrants further exploration to address inherent model limitations and refine evaluation metrics. Moving forward, we aim to enhance the reliability of our methods to better serve real-world applications, aiding researchers in efficiently extracting information and engaging in seamless interactions with clinical documents. Ultimately, these advancements will contribute to enhancing efficiency in both clinical care and research initiatives.
Acknowledgements.
This work was supported by Department of Veterans Affairs, Office of Mental Health and Suicide Prevention. This research used resources of the Knowledge Discovery Infrastructure at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725 and the Department of Veterans Affairs Office of Health Informatics and by VA-DoD Joint Incentive fund under IAA No. 36C10B21M0005. Notice: This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan). The authors wish to acknowledge the support of the larger partnership. Most importantly, the authors would like to thank and acknowledge the veterans who chose to get their care at the VA. Disclaimer: The views and opinions expressed in this manuscript are those of the authors and do not represent those of the Department of Veterans Affairs, the Department of Energy, or the United States Government.